# FedSRD: Sparsify-Reconstruct-Decompose for Communication-Efficient Federated Large Language Models Fine-Tuning

### v2

@@ Abstract.
The current paradigm of training large language models (LLMs) on publicly available Web data is becoming unsustainable, with high-quality data sources in specialized domains nearing exhaustion. Federated Learning (FL) emerges as a practical solution for the next generation of AI on a decentralized Web, enabling privacy-preserving collaborative fine-tuning by leveraging private data distributed across a global client base. While Low-Rank Adaptation (LoRA) is the standard for efficient fine-tuning, its application in federated settings presents a critical challenge: communication overhead remains a significant bottleneck across the Web’s heterogeneous network conditions. The structural redundancy within LoRA parameters not only incurs a heavy communication burden but also introduces conflicts when aggregating client updates. To address this, we propose FedSRD, a Sparsify-Reconstruct-Decompose framework designed for communication-efficient federated LLMs fine-tuning. We first introduce an importance-aware sparsification method that preserves the structural integrity of LoRA updates to reduce the uploaded parameter count. The server then reconstructs and aggregates these updates in a full-rank space to mitigate conflicts. Finally, it decomposes the global update into a sparse low-rank format for broadcast, ensuring a symmetrically efficient cycle. We also propose an efficient variant, FedSRD-e, to reduce computational overhead. Experimental results on 10 benchmarks demonstrate that our framework significantly reduces communication costs by up to 90% while even improving model performance on heterogeneous client data.
Federated Learning, Large Language Models, Communication, LoRA, Fine-tuning
copyright: nonedoi: XXXXXXX.XXXXXXXccs: Computing methodologies Cooperation and coordination
@@ 1. Introduction
Large Language Models (LLMs), pre-trained on vast web-scale corpora, have demonstrated remarkable generalist capabilities (Ouyang et al., 2022; Kaplan et al., 2020). However, realizing their full potential for specialized downstream applications requires fine-tuning on diverse, high-quality, domain-specific data. Such data is often proprietary and distributed across numerous private sources, resulting in fragmented ”data silos”. As centralizing this sensitive data is frequently impractical due to privacy regulations and cost constraints, Federated Learning (FL) offers a compelling solution, enabling multiple clients to collaboratively fine-tune a global shared LLM without direct data exchange. This paradigm is central to building the next generation of intelligent, decentralized Web applications (Sani et al., [n. d.]; Wu et al., 2025).
While FL provides a privacy-preserving solution, fine-tuning billion-parameter LLMs across distributed clients remains computationally and communicatively prohibitive for individual participants. To overcome this, Parameter-Efficient Fine-Tuning (PEFT) methods, particularly Low-Rank Adaptation (LoRA) (Hu et al., 2022), have become the standard (Ding et al., 2023; Kuang et al., 2024; Ye et al., 2024). As illustrated in Figure 1, LoRA significantly reduces the training burden by freezing the pre-trained model weights and optimizing only a small set of trainable, low-rank matrices. In a federated setting, clients train and communicate only these LoRA parameters, achieving performance comparable to full-parameter fine-tuning at a fraction of the resource cost.
Despite the efficiency of LoRA, the iterative communication of its parameters remains a substantial overhead. This becomes a critical bottleneck when deployed across the highly heterogeneous and unpredictable network conditions characteristic of the global Web and mobile systems (Dorfman et al., 2023; Mundt, 2004). High communication costs can exclude clients with limited bandwidth, such as those in rural areas or on mobile devices, undermining data diversity and posing fairness and inclusivity challenges for the global Web ecosystem. While sparsification is a common strategy for reducing communication, conventional methods are ill-adapted for federated LoRA fine-tuning on non-IID data. As shown in Figure 2, naïve approaches like magnitude-based pruning (Aji and Heafield, 2017) disrupt the critical structural relationship between the LoRA matrices $A$ and $B$, while random sparsification (Yu et al., 2024) also leads to performance degradation at high sparsity ratios. Crucially, both methods fail to preserve client-specific knowledge and exacerbate conflicts during server-side aggregation under data heterogeneity. This results in an undesirable trade-off: reduced communication at the expense of significant performance loss. This leads us to our central research question:
@@
To address this question, we introduce FedSRD, a novel Sparsify-Reconstruct-Decompose framework for communication-efficient federated LLMs fine-tuning. Our framework is designed to maintain high model performance through three sequential operations.
First, on the client side, we develop an importance-aware Sparsification strategy. Instead of relying on parameter magnitude, we calculate the importance scores that quantify the contribution of each element in the update matrices ($\Delta B$ and $\Delta A$) to the overall low-rank update ($\Delta W$). This effectively preserves the critical structural interplay between the LoRA matrices. Furthermore, recognizing that different layers in an LLM serve distinct functions, FedSRD adaptively determines the sparsity ratio for each layer based on the kurtosis of its importance score distribution. Aggressive pruning is applied where high importance is concentrated in a few parameters, while conservative pruning is adopted where importance is more diffuse. Second, the server executes the Reconstruction and Decomposition. Upon receiving sparse updates, the server first reconstructs each client’s full-rank weight update ($\Delta W$) before aggregation. This enables aggregation to occur in the full-rank parameter space, a process that is mathematically sound and effectively mitigates the update conflicts arising from non-IID data (Wang et al., 2024; Sun et al., 2024). Finally, to create a symmetrically efficient communication cycle, the server performs a novel Decomposition of the aggregated global update back into a sparse, low-rank format for clients to download. Using a Taylor expansion-based approximation, we employ an alternating scheme that solves for and transmits only one sparse low-rank matrix per round. Furthermore, for scenarios where server-side computation is the bottleneck (e.g., on edge servers), we propose an efficient variant, FedSRD-e, which bypasses the computationally intensive SVD projection during decomposition.
We demonstrate the effectiveness of our framework through extensive experiments that simulate practical non-IID scenarios by finetuning an LLM with clients from diverse specialist domains. Our evaluation on 8 in-domain and 2 out-of-domain benchmarks shows that FedSRD reduces both upload and download communication by 90% while achieving even better performance. In summary, our contributions are:
@@ 2.1. Parameter-Efficient Fine-Tuning
The large scale of modern LLMs makes the full-parameter fine-tuning computationally prohibitive in many resource-constrained environments, such as on mobile or edge devices. To address this, Parameter-Efficient Fine-Tuning (PEFT) methods (Ding et al., 2023) have been developed, which freeze most of a pre-trained model’s parameters and only update a small part. Current approaches include adapter-based methods (He et al., 2021; Edalati et al., 2025; Houlsby et al., 2019; Hu et al., 2022) that insert small, trainable modules between model layers, and prompting methods (Petrov et al., 2023; Lester et al., 2021), such as prefix-tuning (Li and Liang, 2021), that add learnable tokens to the input sequence. Among these, Low-Rank Adaptation (LoRA) (Hu et al., 2022) has emerged as a dominant and effective approach. The key insight of LoRA is that weight updates during fine-tuning possess a low intrinsic rank. LoRA represents the update for each target weight matrix as the product of two smaller, trainable low-rank matrices, significantly reducing the number of trainable parameters and associated memory footprint while achieving comparable performance.
@@ 2.2. Federated LLMs Fine-tuning
Federated Learning (FL) enables the collaborative fine-tuning of LLMs on decentralized private data, addressing both data scarcity and privacy concerns. A common approach is to combine PEFT methods with standard aggregation algorithms (Long et al., 2024; Kuang et al., 2024; Zhang et al., 2024; Nguyen et al., 2024; Ye et al., 2024). Recent studies point out that this naive combination introduces aggregation inaccuracy, where separately averaging the decomposed LoRA matrices ($B$ and $A$) from clients is not mathematically equivalent to averaging the true underlying model updates ($BA$). Several advanced methods have been proposed to mitigate this issue. FFA-LoRA (Sun et al., 2024) ensures a mathematically sound aggregation by freezing the LoRA $A$ matrices and only aggregating the $B$ matrices. FLoRA (Wang et al., 2024) stacks clients’ LoRA matrices to ensure accurate aggregation. FlexLoRA (Bai et al., 2024) proposes to aggregate the true model updates ($BA$) on the server and then employs Singular Value Decomposition (SVD) to derive new global low-rank matrices for clients. FedICU (Liao et al., [n. d.]) disentangles client updates into consensus and divergence components, which are then aggregated separately.
@@ 4. Proposed Method
In this section, we introduce FedSRD, a Sparsify-Reconstruct-Decompose framework for federated LoRA fine-tuning, illustrated in Figure 3. FedSRD establishes a symmetrically efficient communication cycle by transmitting only sparse updates of the LoRA parameters. At each communication round $t$, a client $i$ first computes its local LoRA updates $\Delta B_{i}^{t}$ and $\Delta A_{i}^{t}$, then uses our proposed importance-aware strategy to sparsify them for upload (Section 4.1). The server then reconstructs each client’s full weight matrix and aggregates them in the full-rank space to mitigate non-IID conflicts (Section 4.2). Finally, to ensure an efficient download, the server decomposes the resulting global update into a single sparse low-rank matrix for broadcast using an alternating, Taylor approximation-based method (Section 4.3). The pseudo-code is provided in Appendix A.
@@ 4.1. Importance-aware Sparsification
A higher importance score indicates a greater contribution to the final weight update. We explain the rationale behind in Appendix B. We then prune elements with the lowest scores, transmitting only the most important portion of the parameters.
Furthermore, recognizing that LoRA matrices in different layers play distinct roles (Luo et al., 2025), we propose to adaptively assign a sparsity ratio $\rho$ to each matrix. The core intuition is to use the kurtosis of the importance score distribution reveals a layer’s functional ”prunability”. A high kurtosis (a leptokurtic, heavy-tailed distribution) signifies that importance is concentrated in a few parameters, justifying an aggressive sparsity ratio. Conversely, a low kurtosis (a platykurtic, flat distribution) implies that importance is diffused, necessitating a more conservative ratio. The layer-specific sparsity ratio is thus calculated as:
@@
where $\alpha$ is a hyperparameter. The sparsity ratios are truncated if they exceed a pre-set upper bound to ensure validity. We provide a detailed explanation and justification of this design in Appendix F. We use the adaptive sparsity ratios to prune $\Delta B_{i}^{t}$ and $\Delta A_{i}^{t}$:
@@ 4.2. Reconstruct to Aggregation
Most existing federated learning methods like FedAvg (McMahan et al., 2017), when applied to LoRA, average the clients’ LoRA matrices independently. This aggregation is not mathematically sound and introduces unexpected noises because it multiplicative relationship between the matrices and fails to mitigate update conflicts arising from non-IID client data (Sun et al., 2024; Bai et al., 2024; Wang et al., 2024). This motivates us to develop an advanced aggregation method.
@@ 4.3. Alternative Decomposition
where the higher-order term $\Delta B^{t}\Delta A^{t}$ is omitted (discussed in Appendix G). To solve this underdetermined system, we alternate between setting one of the deltas to zero. In even-numbered rounds, we set $\Delta B^{t}=\mathbf{0}$ and solve for $\Delta A^{t}$:
@@ 5.1.2. Baselines
We compare our framework FedSRD and FedSRD-e with 6 state-of-the-art federated learning methods: (1) FedAvg (McMahan et al., 2017), (2) FedProx (Li et al., 2020), (3) FFA-LoRA (Sun et al., 2024), which freezes LoRA $A$ matrcies to ensure efficient communication and effective aggregation, (4) FlexLoRA (Bai et al., 2024), which aggregates in full-rank space and uses SVD for download, (5) FedICU (Liao et al., [n. d.]), which selectively updates on global model and perform consensus-divergence splitting in the server-side aggregatio, (6) Fed-DARE, which applies random sparsification with rescaling (DARE (Yu et al., 2024)) to LoRA updates for upload. A detailed description of baselines is in Appendix C.
@@ 5.1.4. Implementation
Table 3. Comparison of model performance and per-round communication cost using Llama3.2-3B as the base model. OOD represents Out-of-Domain. The best and second-best results are highlighted in bold and underline, respectively.
Method
Comm.
Cost
Average In-Domain Out-of-Domain
@@
Table 4. Comparison of model performance and per-round communication cost using Qwen2-7B as the base model. OOD represents Out-of-Domain. The best and second-best results are highlighted in bold and underline, respectively.
Method
Comm.
Cost
Average In-Domain Out-of-Domain
In-Domain OOD Code Medical Finance Math AGIEval Law
HumanEval MBPP MedQA MedMCQA FinEval FinanceIQ GSM8K MATH
@@ 5.2.1. Performance Comparison
The main results, reported in Table 3 (Llama3.2-3B) and Table 4 (Qwen2-7B), comprehensively demonstrate the effectiveness and efficiency of our FedSRD framework. We analyze performance from three key perspectives.
@@
Across the experiments, our proposed methods, FedSRD and FedSRD-e, consistently achieve state-of-the-art performance on in-domain benchmarks. As shown in Table 3, with the Llama3.2-3B model, FedSRD and FedSRD-e achieve the highest average in-domain performance score, outperforming all baseline methods. This trend is further confirmed with the larger Qwen2-7B model in Table 4, where FedSRD and FedSRD-e again lead with top performance, respectively. While some baselines may achieve peak performance on specific benchmarks, they often falter in other domains. In contrast, our framework exhibits robust and consistently high performance across all four specialized domains, underscoring its superior capability in aggregating diverse knowledge from specialist clients to build powerful global LLMs.
@@
A standout advantage of our framework is its remarkable communication efficiency. FedSRD and FedSRD-e reduce the per-round communication overhead to just 74 MB for Llama3.2-3B and 61 MB for Qwen2-7B. This represents a reduction of $\approx$ 90% compared to standard methods like FedAvg (742 MB) and advanced methods like FlexLoRA (742 MB). This reduction is achieved by our proposed importance-aware sparsification for uploads and alternative decomposition for downloads (see Appendix D for a detailed analysis). Crucially, this significant gain in efficiency does not come at the cost of performance. On the contrary, FedSRD and FedSRD-e achieve higher performance than baselines, directly answering our central research question. This is visualized in Figures 4 and 5, where our methods are shown to be Pareto optimal: achieving the best performance with the least communication.
@@
Our methods also demonstrate strong performance on out-of-domain (OOD) benchmarks, indicating effective mitigation of catastrophic forgetting. For Llama3.2-3B, FedSRD and FedSRD-e achieve the leading OOD performance. For Qwen2-7B, they also maintain competitive performance. Although FFA-LoRA also exhibits strong OOD retention, we attribute this to its mechanism of freezing the LoRA $A$ matrices. which inherently constrains the capacity and hinders new knowledge acquisition in favor of preserving prior information. In contrast, our methods provide a more balanced and consistently high level of performance across both in-domain and out-of-domain evaluations. Further, we posit that the SVD projection in FedSRD, by seeking a low-rank consensus for in-domain tasks, may inadvertently filter out some general knowledge from the base model, which could account for the superior OOD performance of FedSRD-e. In summary, our framework not only excels at integrating specialized knowledge but also effectively preserves the general capabilities of the base model.

### v3

@@ Abstract.
The current paradigm of training large language models (LLMs) on public available Web data is becoming unsustainable as high-quality data sources in specialized domains near exhaustion. Federated Learning (FL) emerges as a practical solution for the next generation of AI on a decentralized Web, enabling privacy-preserving collaborative fine-tuning on decentralized private data. While Low-Rank Adaptation (LoRA) is standard for efficient fine-tuning, its federated application faces a critical bottleneck: communication overhead under heterogeneous network conditions. Structural redundancy in LoRA parameters increases communication costs and causes aggregation conflicts. To address this, we propose FedSRD, a Sparsify-Reconstruct-Decompose framework for communication-efficient federated LLM fine-tuning. We introduce importance-aware sparsification to reduce the upload parameter count while preserving the structural integrity of LoRA updates. The server aggregates updates in full-rank space to mitigate conflicts, then decomposes the global update into a sparse low-rank format for broadcast, ensuring a symmetrically efficient cycle. We also propose an efficient variant, FedSRD-e, to reduce computational overhead. Experiments on 10 benchmarks show our framework significantly reduces communication costs by up to 90% while improving performance on heterogeneous client data.
Federated Learning; Large Language Models; Communication; LoRA
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/3774904.3792144ccs: Computing methodologies Cooperation and coordination
@@ 1. Introduction
Large Language Models (LLMs), pre-trained on vast web-scale corpora, have demonstrated remarkable generalist capabilities (Ouyang et al., 2022; Kaplan et al., 2020). However, realizing their full potential for specialized downstream applications requires fine-tuning on diverse, high-quality, domain-specific data. Such data is often proprietary and distributed across numerous private sources, resulting in fragmented ”data silos”. As centralizing this sensitive data is frequently impractical due to privacy regulations and cost constraints, Federated Learning (FL) offers a compelling solution, enabling multiple clients to collaboratively fine-tune a global shared LLM without direct data exchange. This paradigm is central to building the next generation of intelligent, decentralized Web applications (Sani et al., ; Wu et al., 2025).
While FL provides a privacy-preserving solution, fine-tuning billion-parameter LLMs across distributed clients remains computationally and communicatively prohibitive. To overcome this, Parameter-Efficient Fine-Tuning (PEFT) methods, particularly Low-Rank Adaptation (LoRA) (Hu et al., 2022), have become the standard (Ding et al., 2023; Kuang et al., 2024; Ye et al., 2024). As illustrated in Figure 1, LoRA significantly reduces the training burden by optimizing only a small set of trainable, low-rank matrices. In a federated setting, clients train and communicate only these LoRA parameters, achieving performance comparable to full-parameter fine-tuning at a fraction of the resource cost.
Despite the efficiency, the iterative communication of LoRA parameters remains a substantial overhead. This becomes a critical bottleneck when deployed across the heterogeneous and unpredictable network conditions characteristic of the global Web and mobile systems (Dorfman et al., 2023; Mundt, 2004). High communication costs can exclude clients with limited bandwidth, such as those in rural areas or on mobile devices, undermining data diversity and posing fairness and inclusivity challenges for the global Web ecosystem. While sparsification is a common strategy for reducing communication, conventional methods are ill-adapted for federated LoRA fine-tuning on non-IID data. As shown in Figure 2, naïve approaches like magnitude-based pruning (Aji and Heafield, 2017) disrupt the structural relationship between LoRA matrices $A$ and $B$, while random sparsification (Yu et al., 2024) also degrades at high sparsity ratios. Crucially, both methods fail to preserve client-specific knowledge and exacerbate conflicts during server-side aggregation under data heterogeneity. This results in an undesirable trade-off: reduced communication at the expense of significant performance loss. This leads to our central research question:
@@
To address this question, we introduce FedSRD, a novel Sparsify-Reconstruct-Decompose framework for communication-efficient federated LLMs fine-tuning. Our framework ensures high model performance through three sequential operations.
First, on the client side, we develop an importance-aware Sparsification strategy. Instead of relying on parameter magnitude, we calculate the importance scores that quantify the contribution of each element in update matrices ($\Delta B$ and $\Delta A$) to the overall update ($\Delta W$). This effectively preserves the critical structural interplay between LoRA matrices. Furthermore, recognizing that different layers serve distinct functions, we adaptively determine layer-wise sparsity ratios based on the kurtosis of its importance scores. Aggressive pruning is applied where high importance is concentrated. Second, the server executes the Reconstruction and Decomposition. Upon receiving sparse updates, the server first reconstructs each client’s full-rank update ($\Delta W$) before aggregation. This enables aggregation to occur in the full-rank space, which is mathematically sound and effectively mitigates the update conflicts from non-IID data (Wang et al., 2024; Sun et al., 2024). Finally, to create a symmetrically efficient communication cycle, the server performs a novel Decomposition of the aggregated global update back into a sparse, low-rank format for clients to download. Using a Taylor expansion-based approximation, we employ an alternating scheme that solves for and transmits only one sparse low-rank matrix per round. Furthermore, We propose FedSRD-e, a computational-efficient variant that bypasses intensive SVD operations for resource-constrained servers.
We demonstrate the effectiveness of FedSRD through extensive experiments that simulate practical non-IID scenarios by finetuning an LLM with clients from diverse specialist domains. Our evaluation on 8 in-domain and 2 out-of-domain benchmarks shows that FedSRD reduces both upload and download communication by 90% while achieving even better performance. In summary, our contributions are:
@@ 2.1. Parameter-Efficient Fine-Tuning
The large scale of modern LLMs makes the full-parameter fine-tuning computationally prohibitive in resource-constrained environments. To address this, Parameter-Efficient Fine-Tuning (PEFT) methods (Ding et al., 2023) freeze most of model parameters and only update a small part. Current approaches include adapter-based methods (He et al., 2021; Edalati et al., 2025; Houlsby et al., 2019; Hu et al., 2022) that insert small, trainable modules in models, and prompting methods (Petrov et al., 2023; Lester et al., 2021), such as prefix-tuning (Li and Liang, 2021), that add learnable tokens to the input sequence. Among these, Low-Rank Adaptation (LoRA) (Hu et al., 2022) has emerged as a dominant and effective approach. The key insight of LoRA is that weight updates in fine-tuning possess a low intrinsic rank. LoRA represents each weight matrix update as the product of two smaller, trainable low-rank matrices, significantly reducing the number of trainable parameters and memory footprint while achieving comparable performance.
@@ 2.2. Federated LLMs Fine-tuning
Federated Learning (FL) enables the collaborative fine-tuning of LLMs on decentralized private data, addressing both data scarcity and privacy concerns. A common approach is to combine PEFT methods with standard aggregation algorithms (Long et al., 2024; Kuang et al., 2024; Zhang et al., 2024; Nguyen et al., 2024; Ye et al., 2024). Recent studies point out that this naive combination introduces aggregation inaccuracy, where separately averaging the decomposed LoRA matrices ($B$ and $A$) from clients is not mathematically equivalent to averaging the true underlying model updates ($BA$). Several advanced methods have been proposed to mitigate this issue. FFA-LoRA (Sun et al., 2024) ensures a mathematically sound aggregation by freezing the LoRA $A$ matrices and only aggregating the $B$ matrices. FLoRA (Wang et al., 2024) stacks clients’ LoRA matrices to ensure accurate aggregation. FlexLoRA (Bai et al., 2024) proposes to aggregate the true model updates ($BA$) on the server and then employs Singular Value Decomposition (SVD) to derive new global low-rank matrices for clients. FedICU (Liao et al., ) disentangles client updates into consensus and divergence components, which are then aggregated separately.
@@ 4. Proposed Method
In this section, we introduce FedSRD, a noval Sparsify-Reconstruct-Decompose framework for federated LoRA fine-tuning, illustrated in Figure 3. At each communication round $t$, a client $i$ first computes its local LoRA updates $\Delta B_{i}^{t}$ and $\Delta A_{i}^{t}$, then uses our proposed importance-aware strategy to sparsify them for upload (Section 4.1). The server then reconstructs each client’s full weight matrix and aggregates them in the full-rank space to mitigate non-IID conflicts (Section 4.2). Finally, to ensure an efficient download, the server decomposes the resulting global update into a single sparse low-rank matrix for broadcast using an alternating, Taylor approximation-based method (Section 4.3).
@@ 4.1. Importance-aware Sparsification
A higher importance score indicates a greater contribution to the final weight update. We then prune elements with the lowest scores. We explain the rationale behind and prove that pruning based on this score minimizes the upper bound of the weight update approximation error $\|\Delta W_{i}^{t}\|_{F}$ in Appendix B.
Furthermore, recognizing that LoRA matrices in different layers play distinct roles (Luo et al., 2025), we adaptively assign a sparsity ratio $\rho$ to each matrix. The core intuition is that the kurtosis of the importance score reveals a layer’s functional ”prunability”. A high kurtosis (a leptokurtic, heavy-tailed distribution) signifies that importance is concentrated in a few parameters, justifying an aggressive sparsity ratio. Conversely, a low kurtosis (a platykurtic, flat distribution) implies that importance is diffused, necessitating a more conservative ratio. The layer-specific sparsity ratio is thus calculated as:
@@
where $\alpha$ is a hyperparameter. The sparsity ratios are truncated if they exceed a pre-set upper bound to ensure validity. We provide a detailed explanation and justification of this design in Appendix E. We use the adaptive sparsity ratios to prune $\Delta B_{i}^{t}$ and $\Delta A_{i}^{t}$:
@@ 4.2. Reconstruct to Aggregation
Most existing federated learning methods like FedAvg (McMahan et al., 2017), when applied to LoRA, average the clients’ LoRA matrices independently. This aggregation is not mathematically sound and introduces unexpected noises because its multiplicative relationship between the matrices and fails to mitigate update conflicts arising from non-IID client data (Sun et al., 2024; Bai et al., 2024; Wang et al., 2024). This motivates us to develop an advanced aggregation method.
@@ 4.3. Alternative Decomposition
where the higher-order term $\Delta B^{t}\Delta A^{t}$ is omitted (discussed in Appendix F). To solve this underdetermined system, we alternate between setting one of the deltas to zero. In even-numbered rounds, we set $\Delta B^{t}=\mathbf{0}$ and solve for $\Delta A^{t}$:
@@ 5.1.2. Baselines
We compare our framework FedSRD and FedSRD-e with 6 state-of-the-art federated learning methods: (1) FedAvg (McMahan et al., 2017), (2) FedProx (Li et al., 2020), (3) FFA-LoRA (Sun et al., 2024), which freezes LoRA $A$ matrcies to ensure efficient communication and effective aggregation, (4) FlexLoRA (Bai et al., 2024), which aggregates in full-rank space and uses SVD for download, (5) FedICU (Liao et al., ), which selectively updates on global model and perform consensus-divergence splitting in the server-side aggregatio, (6) Fed-DARE, which applies random sparsification with rescaling (DARE (Yu et al., 2024)) to LoRA updates for upload.
@@ 5.1.4. Implementation
Table 3. Comparison of model performance and per-round communication cost. The top section uses Llama3.2-3B and the bottom section uses Qwen2-7B. OOD represents Out-of-Domain. Best and second-best results are bolded and underlined.
Method Comm. Cost (MB) Average In-Domain Out-of-Domain
@@
Base Model: Qwen2-7B
@@ 5.2.1. Performance Comparison
The main results are reported in Table 3. We analyze the results from three key perspectives.
@@
Across the experiments, our proposed methods, FedSRD and FedSRD-e, consistently achieve state-of-the-art performance on in-domain benchmarks. As shown in Table 3, with the Llama3.2-3B model, FedSRD and FedSRD-e achieve the highest average in-domain performance score, outperforming all baseline methods. This trend is further confirmed with the larger Qwen2-7B model, where FedSRD and FedSRD-e again lead with top performance, respectively. While some baselines may achieve peak performance on specific benchmarks, they often falter in other domains. In contrast, our framework exhibits robust and consistently high performance across all four specialized domains, underscoring its superior capability in aggregating diverse knowledge from specialist clients to build powerful global LLMs.
@@
A standout advantage of our framework is its remarkable communication efficiency. FedSRD and FedSRD-e reduce the per-round communication overhead to just 74 MB for Llama3.2-3B and 61 MB for Qwen2-7B, an approximate 90% reduction compared to standard methods like FedAvg (742 MB) and advanced methods like FlexLoRA (742 MB). This is achieved by our proposed importance-aware sparsification for uploads and alternative decomposition for downloads (see Appendix C for a detailed analysis). Crucially, this significant gain in efficiency does not compromise performance. On the contrary, FedSRD and FedSRD-e achieve higher performance than baselines, directly answering our central research question. This is visualized in Figures 4 and 5, where our methods are shown to be Pareto optimal: achieving the best performance with the least communication.
@@
Our methods also show strong performance on out-of-domain (OOD) benchmarks, effectively mitigating catastrophic forgetting. FedSRD and FedSRD-e achieve leading results on Llama3.2-3B and remain competitive on Qwen2-7B. Although FFA-LoRA exhibits strong OOD retention, we attribute this to its mechanism of freezing the LoRA $A$ matrices. which inherently constrains the capacity and hinders new knowledge acquisition in favor of preserving prior information. In contrast, our methods provide a more balanced and consistently high performance across both in-domain and out-of-domain evaluations. Further, we posit that the SVD projection in FedSRD, by seeking a low-rank consensus for in-domain tasks, may inadvertently filter out some general knowledge from the base model, which could account for the superior OOD performance of FedSRD-e. In summary, our framework not only excels at integrating specialized knowledge but also effectively preserves the general capabilities of the base model.